how to calculate log to the base of 10 (1.05) without a calculator and the negative log of this number

Grade:12th pass

1 Answers

Deepak Kumar Shringi
5 years ago
for this you should remember some basic values
Log Base 10 of... Is equal to...
2 0.3010
3 0.4771
5 0.698
6 0.778
7 0.845
8 0.903
log1.05=log105/100=log21/20=log21-log20=(log3+log7)/log5+2log2=1.322/1.3=1.017
(also you should be aware of basic properties of log_}